OpenStudy (anonymous):

find limit as x approaches infinity. (5x^2+6)/(6x^2)

OpenStudy (anonymous):

\[\frac{5}{6}\] with your eyeballs

OpenStudy (anonymous):

same power

OpenStudy (anonymous):

numerator and denominator of are the same degree, so it is ratio of leading coefficients.

OpenStudy (anonymous):

it would factor down to (5+6/x^2)/6 ?

OpenStudy (anonymous):

looks like \[\frac{5}{6}\] if you do it with numbers. put x = 100 you get \[\frac{50000}{60006}\]

OpenStudy (anonymous):

ok so you just canceled out the 6/x^2?

OpenStudy (anonymous):

thats in the numerator